Graph using the intercept method 5x - 10y = 20
Answer by wuwei96815(245)

Let x = 0
Then:
5x-10y=20
5(0)-10y=20
-10y=20
y=-2

Let y = 0
Then:
5x-10y=20
5x-10(0)=20
5x=20
x=4

Now the cordinates (0,-2) and (4,0) can be used to plot your line.

Did you follow all of that?